# Speed-Accuracy Trade-Off in the Simulation of Realistic Arterial Blood Flow: Effects of Geometric Simplifications and Non-Newtonian Models

### v1

@@ 3.3.1 Wall shear stress (WSS)
Wall shear stress (WSS) plays a crucial role in regulating endothelial function and has been widely recognized as a key factor in the development of atherosclerosis. Endothelial cells sense the wall shear stress via multiple responsive cell components, and activation (or) deactivation of intra-cellular pathways leads to gene and protein expression regulation[31]. This continuous process is important for the remodeling, development, and growth of the vascular walls. Moreover, clinical significance of wall shear stress lies in its ability to predict the risk of future cardiovascular events, making it an important biomarker in cardiovascular health. Mathematically, the WSS can be calculated with the help of velocity gradients. $(\nabla\textbf{u})$ as follows:
@@
The strain rate ($\bar{\bar{\gamma}}$) is given by the sum of the velocity gradient and the transpose of the velocity gradient $\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$ at the wall. The magnitude of the strain rate tensor field can be calculated as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}\colon\bar{\bar{\gamma}}}$. The wall shear stress is calculated for the patient-specific and simplified geometries (all levels) and presented in the form of contours.
@@
The wall shear stress (WSS) contours are shown in Figure 10 for all the time instants of the pulsatile cycle for the original and other geometric simplification levels while adopting the Apostolidis and Beris viscosity model. The features of WSS in Level 1 are particularly significant in predicting the onset and progression of disease, as they provide detailed insights, as well as the identification of the critical regions. Figure 10 shows how the simulation accuracy deteriorates by simplification of the geometry. Firstly, for Level 1, with all perturbations preserved, we observe high variability in WSS at the peak instant of flow. At all times, the surface perturbations and the asymmetry generate several high (and low) stress zones of WSS. Some of these variabilities are captured in Level 2, although the asymmetry of Level 1 is lost. The high and low-stress zones appear as bands around the circumference, with an overall reduction of the variations observed in the WSS distributions of Level 1. This is most pronounced at the time instant of peak velocity. However, at Level 3 (tube of decreasing area), the WSS profile is smooth and all the variability is lost. Thus, Level 2 might still provide satisfactory predictions of WSS, which is entirely lost in Level 3.
3.3.2 Time averaged wall shear stress (TAWSS)
@@
Here, $T$ is the time period of the cardiac cycle. Time-averaged wall shear stress (TAWSS) calculated for the fourth cycle of the simulations is shown in Figure 11 for the three levels of the geometry of the descending aorta and selected viscosity models. In Figure 11 (i),(ii) and (iii) TAWSS contours are plotted for Apostolidis and Beris, Carreau Yasuda, and Newtonian viscosity models, respectively. The trends are qualitatively similar for WSS discussed earlier. In Level 1, the perturbations and asymmetry of the geometry generate a significant amount of zones for high and low TAWSS. Some of these zones are retained in Level 2 with less asymmetry and a reduction in the shear stress magnitudes. A smooth variation of TAWSS is obtained in Level 3 which is a smooth geometry. The trends are similar across the three viscosity models. However, the perturbations in TAWSS are slightly different with the Apostolidis and Beris model, whereas the profiles are similar for the Newtonian and Carreau-Yasuda models. The values are slightly higher with the Apostolidis and Beris model in Level 1. Thus, again, concerning spots of extreme stress, Level 2 captures the essence partially while Level 3 shows a smooth variation.
@@ 3.3.3 Oscillatory shear index (OSI)
As discussed earlier, WSS levels activate and deactivate the endothelial cell response. In pulsatile flow, the shear stress near the artery wall fluctuates, making the endothelial cell deliver a pro-inflammatory response. The Oscillatory Shear Index (OSI) introduced by Ku et al. [32] provides a measure of these fluctuations. The OSI can be calculated as follows:
@@
From the definition of OSI, the high-value region shows a high time-periodic oscillation of the wall shear stress. OSI varies between 0 and 0,5, the latter representing a purely oscillatory response.
@@
Figure 12b plots Oscillatory Shear Index Contours for the selected viscosity models. OSI contours depict the zone of shear oscillation and its intensity. In Figure 12, the OSI distributed zones in the Level 1 artery can be seen with hotspots of high and low OSI values, signifying the nature of shear stress oscillations. The region with high OSI values is more prone to a pro-inflammatory response by the endothelial cells, whereas the low value of the OSI zone indicates unidirectional shear stress; therefore, an anti-inflammatory response for the cell is obtained. As the complexity of the descending aorta gradually decreases, the OSI zones are also averaged uniformly as the geometrical features are toned down in the subsequent levels of simplification. Lastly, a uniform OSI distribution is observed in Level 3, as no geometrical disturbances are present. It is clear that the asymmetry of Level 1 is lost in both Levels 2 and 3 due to the geometric simplification. Thus, a fair comparison of the time-averaged hemodynamic indicators across geometric levels is shown in Fig. 13 for TAWSS (Fig. 13 (a)) and OSI (Fig.13 (b)) averaged over the circumference at each axial location. As might be expected, the predictions in Level 3 are a monotonic function, while predictions in Levels 1 and 2 show oscillations due to axial variations in the geometry. For all data shown, the Level 3 predictions represent an average behavior of the oscillatory predictions obtained for Levels 1 and 2. The usefulness of Level 2 is readily visible in these results since, mostly, it closely follows the predictions of Level 1. In the present discussion, the Level 1 predictions are averaged over the circumference and the Level 2 geometry is built by averaging diameters at various axial locations. The consistency of Levels 1 and 2 is visible even more for the non-Newtonian models and shows some differences for the Newtonian fluid, especially for OSI. These results bolster the proposition that a somewhat smoother geometry (like Level 2) will be able to capture the fluctuations in hemodynamic parameters, although the local asymmetry will be lost. As discussed later in Section 3.5, this presents a valid ground for a speed-accuracy trade-off since convergence in the smoother geometries is faster and yet captures the essence of the predictions of Level 1. However, only Level 1 (patient-specific geometry) can be used without any further simplifications in applications where the details of local asymmetry are required.
@@
For TAWSS, Level 2 is able to provide predictions close to those of the circumferential averaged Level 1 data. However, Level 3 is only able to show a monotonic function representing an overall average trend. It lacks the oscillation and fluctuations that are caused due to area fluctuation, characteristic of Levels 1 and 2. It is possible to ask, from the prediction of TAWSS of Level 3 and knowing the area changes, is it possible to recover the prediction of Levels 1 and 2? The following approach is adopted. Shear stress is given by
@@
Here, symbols $U$ and $D$ are the characteristic velocity and diameter at any location within the duct. The average velocity at any location can be estimated by knowing the volumetric flow rate and local cross-sectional area. Since we know the total volumetric flow rate and area at any location, $U$ and $D$ can be estimated for local perturbations while $\frac{u*}{z*}$ can be assumed to be unaffected. At a high shear rate, $\mu_{app}$ can be assumed to be a limiting value (Figure 3). Thus, TAWSS can be estimated locally by estimating the characteristic velocity and diameter at every cross-section. These estimates are shown in Figure 13(a) for all viscosity models by black dashed lines.
@@ 3.4 $Q$ criterion
Another critical parameter for studying strain rate and rotational characteristics is the variation of the $Q$ criterion. $Q$ criterion is based on the velocity gradient tensor, which is given as the summation of a symmetric and anti-symmetric part as follows:
@@
A positive $Q$ value shows the region with a dominance of rotational flow. Similarly, a negative value denotes a zone where rotational flows are not dominant. In Figure 14a, an iso-surface of the $Q$-criterion is shown for $Q$ = 0.1. Similarly, in Figure 14b, it is shown for $Q$ = -0.1. It is clear that the simplification of the artery leads to a uniform flow distribution. In Level 1, the $Q$ criterion shows a non-uniform asymmetric distribution; as the geometry is simplified to Level 2, the distribution becomes symmetric and breaks into zones. In Level 3, only a smooth distribution is obtained. In this discussion, the value of the $Q$ function is selected in such a way as to fill the descending aorta. It is observed that in the systole phase, the vortical structure for Level 1 is distributed in all regions. However, for Level 2, the same structure breaks and spreads towards the periphery of the wall. For Level 3, the vortical structures are uniformly distributed over the entire flow space. As the flow rate increases, the structure becomes stronger. However, In the diastolic phase, the strain rate grows and dominates across all levels due to a lower flow rate. Further, separation and breakage of the flow structures are also observed. A similar trend is also seen for the contours of $Q$ = -0.1, as shown in Figure 14b. The nature of the distribution of the $Q$ function is associated with wall shear stress and, correlates closely with those of the hemodynamic indicators, namely, TAWSS and OSI.
@@ 3.5 Speed Accuracy trade-off
Figure 15: Speed-accuracy analysis is carried out by performing (a) grid independence test for simplified geometries up to 25 slices and comparing their respective friction factors. (b) The maximum number of iterations required to reach the tolerance of 1e-5 in velocity is plotted as a function of the element size. (c) Maximum number of iterations is plotted for the Apostolidis and Beris and Newtonian viscosity models for Level 2 (with 25 slices) and Level 3 simplification.
Further, to show the speed increase with geometric simplification, we show the maximum number of iterations required to reach a prescribed convergence criterion as a function of the element size, in Figure 15(b). The data reveal that simplified geometries with 5 to 20 slices require approximately 150 iterations. In contrast, the simplified artery with 25 slices and the geometry with zero slices need around 300 iterations for the smallest element size and 200 iterations for the largest element size. For Level 1 geometry, the number of iterations increases as the element size is decreased, which is consistent with expectations for unstructured meshes over complex three-dimensional geometries. Thus, there is a clear trend of improvement in terms of the number of iterations (and thus, the speed of computations) as we move towards simplified geometric representations. The difference is as large as four times between Level 1 and Level 2 (with 20 or fewer slices) at their optimal element sizes.
In Figure 15(c), the Apostolidis and Beris viscosity model is compared with the Newtonian viscosity model with respect to the maximum number of iterations as a function of element size for Level 2 and Level 3 geometries. A significant difference is observed when the viscosity model becomes non-Newtonian. For the converged mesh, nearly 5000 iterations are required for the non-Newtonian viscosity model as compared to the Newtonian. Overall, geometric simplification reduces the number of iterations required, nonlinearity of the viscosity model increases the number of iterations, while grid refinement increases the number of iterations, the discussion being carried out on meshes with acceptable grid quality.
@@ 5 Conclusions
A detailed study on a patient-specific geometry of the descending aorta with two simplified representations and three viscosity models is conducted. The geometry is extracted from CT scan data shared by a hospital and reconstructed using image processing tools. Realistic time-dependent boundary conditions are applied at the inlet and outlet in terms of pulsatile velocity and outlet pressure, respectively. For the geometrical approximation, the original artery (termed Level 1) is further simplified into two more levels, denoted as Level 2 and Level 3. Level 2 is based on the cross-sections of Level 1, which is divided into 25 sections, with each part having a representative average diameter. Thus, Level 2 captures some of the undulations and area fluctuations of Level 1 but loses the local asymmetry of the original. Level 3 represents a smooth duct based on the inlet and outlet cross-sectional area of the actual artery. The calculations are performed with three blood rheology models – Newtonian, Carreau-Yasuda, and Apostolidis and Beris. Various aspects of blood flow for all the levels of the descending aorta geometry (with selected blood viscosity models) are studied. The following conclusions are arrived at in the present work.
@@
  4. 4.
    A simple estimate of the actual WSS in the original artery can be obtained by knowing the prediction from Level 3, scaled by the area fluctuation. The estimates thus retrieved are pretty close to the ones obtained with Level 1.

### v3

@@ 3.3.1 Wall shear stress (WSS)
Wall shear stress (WSS) plays a crucial role in regulating the endothelial function and has been widely recognized as a key factor in the development of atherosclerosis. Endothelial cells sense the wall shear stress via multiple responsive cell components, and activation or deactivation of intra-cellular pathways leads to gene and protein expression regulation[33]. This continuous process is important for remodeling, development, and growth of the vascular walls. Moreover, clinical significance of wall shear stress lies in its ability to predict the risk of future cardiovascular events, making it an important biomarker in cardiovascular health. Mathematically, WSS can be calculated with the help of velocity gradients $(\nabla\textbf{u})$ as follows:
@@
The strain rate ($\bar{\bar{\gamma}}$) is given by the sum of the velocity gradient and the transpose of the velocity gradient $\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$ at the wall. The magnitude of the strain rate tensor field can be calculated as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}\colon\bar{\bar{\gamma}}}$. Wall shear stress is calculated for the patient-specific and simplified geometries (all levels) and presented in the form of contours. The wall shear stress (WSS) contours are shown in Figure 10 for all the time instants of the pulsatile cycle for the original and other geometric simplification levels while adopting the Apostolidis and Beris viscosity model. The features of WSS in Level 1 are particularly significant in predicting the onset and progression of disease, as they provide detailed insights, as well as the identification of the critical regions.
@@
Figure 10 shows how the simulation accuracy deteriorates by simplifying the geometry. Firstly, for Level 1, with all perturbations preserved, we observe high variability in WSS at the peak instant of flow. At all times, the surface perturbations and the asymmetry generate several high (and low) stress zones of WSS. Some of these variabilities are captured in Level 2, although the asymmetry of Level 1 is lost. The high and low-stress zones appear as bands around the circumference, with an overall reduction of the variations observed in the WSS distributions of Level 1. This is most pronounced at the time instant of peak velocity. However, the WSS profile is smooth at Level 3 (tube of decreasing area), and all the variability is lost. Thus, Level 2 might still provide satisfactory predictions of WSS, which is entirely lost in Level 3. The variability in these hemodynamic parameters, in terms of the maximum and minimum values, are shown in Table 4 in a tabular form. For simplicity, for all three viscosity models, we show the data only at two time instants of the cycle - t = 0 s and t = 0.35 s (peak). The table further highlights the discussion about the accuracy of Levels 2 and 3 (relative to level 1) presented above. The table further shows that while the values in Level 2 retain some similarity, the values for Level 3 are significantly different in most cases. These trends broadly fit the discussion presented above.
Table 4: Minimum and maximum wall shear stress magnitude values recorded at the start (t = 0 s) and at the peak (t = 0.35 s) of the cardiac cycle for the simplified geometrical levels and tabulated for the selected viscosity models.
time Geometry $\rightarrow$ Level 1 Level 2 Level 3
Viscosity models $\downarrow$ min max min max min max
t = 0 s Apostolidis & Beris 0.170 0.647 0.209 0.686 0.230 0.527
Carreau-Yasuda 0.014 0.398 0.037 0.371 0.067 0.229
Newtonian 0.011 0.401 0.394 1.385 0.062 0.223
t = 0.35 s Apostolidis & Beris 0.029 5.930 0.004 3.960 0.378 1.444
Carreau-Yasuda 0.006 2.642 0.000 2.807 0.100 0.764
Newtonian 0.004 2.642 0.001 2.949 0.094 0.756
3.3.2 Correlations
The previous discussion on retention of accuracy is qualitative. To enable the quantification of the impact of simplified geometric levels on accuracy, we introduce a correlation metric here. These correlations can be used to compare the accuracy levels of any hemodynamic parameter across the descending aorta. Here, we show the definition of the correlation for the WSS, but it can also be extended to pressure. The calculations use the deviations of any parameter from its mean at any location, as follows:
[…]
$C_{11}=\frac{\sum_{z}\theta_{1}\theta_{1}}{\sum_{z}\theta_{1}\theta_{1}}=1;C_{12}=\frac{\sum_{z}\theta_{1}\theta_{2}}{\sum_{z}\theta_{1}\theta_{1}};C_{13}=\frac{\sum_{z}\theta_{1}\theta_{3}}{\sum_{z}\theta_{1}\theta_{1}};C_{23}=\frac{\sum_{z}\theta_{2}\theta_{3}}{\sum_{z}\theta_{2}\theta_{2}};$ (11)
where $C_{12}$, $C_{13}$ and $C_{23}$ are the correlation parameter between levels 1 and 2, levels 1 and 3 and levels 2 and 3, respectively. By definition, $C_{11}$ should be unity and is used in the plots for validation. Thus, correlation parameters provide a sense of a “hotspot” (in WSS or pressure) being observed at a similar location when two different geometry levels are used. If a simplified geometry is able to capture all hotspots of level 1 at their respective locations, the value of the correlation parameter would be close to unity.
Here, we evaluate correlation levels for WSS and pressure. These correlations are computed at specific time instants of the waveform. The correlation coefficient, $C$, is evaluated for the wall shear stress (WSS) magnitude and wall pressure over the entire cardiac cycle for the selected viscosity models, and the results are presented in Figure 11. The correlation trends for both WSS (Figure 11(a)) and wall pressure (Figure 11(b)) are examined. The WSS values of $C_{12}$ remain close to unity for the non-Newtonian models but show large fluctuations for the Newtonian model. In contrast, $C_{13}$ and $C_{23}$ remain further away from unity. This quantitatively supports our preceding discussion. For WSS, Level 2 predictions remain highly correlated with Level 1 (and thus retain accuracy), whereas much of this accuracy is lost in Level 3. However, the Newtonian model is not well behaved for these simulations, showing large fluctuations in correlation values across models. Particularly, cross-correlations $C_{12}$ and $C_{13}$ are large during phases C and D for the Newtonian simulation, suggesting the role of the flow structures appearing in Level 1. This perhaps indicates that the non-Newtonian models of blood should be used for such simulations in geometries that contain a high degree of irregularity. Conversely, accuracy loss is observed to be mild in pressure values.
Figure 11: Correlations of (a) wall shear stress magnitude and (b) wall pressure are plotted with time for the selected viscosity model, i.e, (i) Apostilidis and Beris, (ii) Carreau-Yasuda and (iii) Newtonian. The correlation parameter values are highlighted at phases of the velocity waveform given at t = 0 s (A), t = 0.35 s (B), t = 0.55 s (C), and t = 0.9 s (D).
3.3.3 Time-averaged wall shear stress (TAWSS)
@@
Here, $T$ is the time period of the cardiac cycle. Time-averaged wall shear stress (TAWSS) calculated for the fourth cycle of the simulations is shown in Figure 12 for the three levels of the geometry of the descending aorta and selected viscosity models. In Figure 12 (i),(ii) and (iii) TAWSS contours are plotted for Apostolidis and Beris, Carreau Yasuda, and Newtonian viscosity models, respectively. The trends are qualitatively similar to WSS discussed earlier. In Level 1, the perturbations and asymmetry of the geometry generate a significant amount of zones for high and low TAWSS. Some of these zones are retained in Level 2 with less asymmetry and a reduction in the shear stress magnitudes. A gradual variation of TAWSS is obtained in Level 3 which is a smooth geometry. The trends are similar across the three viscosity models. However, the perturbations in TAWSS are slightly different with the Apostolidis and Beris model, whereas the profiles are similar for the Newtonian and Carreau-Yasuda models. The values are slightly higher with the Apostolidis and Beris model in Level 1. Thus, again, concerning spots of extreme stress, Level 2 captures the essence partially while variations in Level 3 are averaged out.
@@ 3.3.4 Oscillatory shear index (OSI)
In pulsatile flow, the shear stress near the artery wall fluctuates, making the endothelial cell deliver a pro-inflammatory response. The Oscillatory Shear Index (OSI) introduced by Ku et al. [34] provides a measure of these fluctuations. The OSI can be calculated as follows:
@@
OSI varies between 0 and 0.5, the latter representing a purely oscillatory response. The high-value region shows a high time-periodic oscillation of the wall shear stress.
@@
Figure 13 plots Oscillatory Shear Index contours for the selected viscosity models. OSI contours depict the zone of shear oscillation and its intensity. In Figure 13, the OSI distributed zones in the Level 1 artery can be seen with hotspots of high and low OSI values, signifying the nature of shear stress oscillations. The region with high OSI values is more prone to a pro-inflammatory response by the endothelial cells, whereas the low value of the OSI zone indicates unidirectional shear stress; therefore, an anti-inflammatory response for the cell is obtained. As the complexity of the descending aorta gradually decreases, the OSI zones are also averaged uniformly as the geometrical features are toned down in the subsequent levels of simplification. Lastly, a uniform OSI distribution is observed in Level 3, as no geometrical disturbances are present. It is clear that the asymmetry of Level 1 is lost in both Levels 2 and 3 due to the geometric simplification.
A fair comparison of the time-averaged hemodynamic indicators across geometric levels is shown in Fig. 14 for TAWSS (Fig. 14 (a)) and OSI (Fig.14 (b)) averaged over the circumference at each axial location. As might be expected, the predictions in Level 3 are a monotonic function, while predictions in Levels 1 and 2 show oscillations due to axial variations in the geometry. For all data shown, the Level 3 predictions represent an average behavior of the oscillatory predictions obtained for Levels 1 and 2. The usefulness of Level 2 is readily visible in these results since, mostly, it closely follows the predictions of Level 1. In the present discussion, the Level 1 predictions are averaged over the circumference and the Level 2 geometry is built by averaging diameters at various axial locations. The consistency of Levels 1 and 2 is visible even more for the non-Newtonian models and shows some differences for the Newtonian fluid, especially for OSI. These results bolster the proposition that a somewhat smoother geometry (like Level 2) will be able to capture the fluctuations in hemodynamic parameters, although the local asymmetry will be lost. As discussed later in Section 3.5, this presents a valid ground for a speed-accuracy trade-off since convergence in the smoother geometries is faster and yet captures the essence of the predictions of Level 1. However, only Level 1 (patient-specific geometry) can be used without any further simplifications in applications where the details of local asymmetry are required.
@@
For TAWSS, Level 2 is able to provide predictions close to those of the circumferential averaged Level 1 data. However, Level 3 is only able to show a monotonic function representing an overall average trend. It lacks the oscillation and fluctuations that are caused due to area fluctuation, characteristic of Levels 1 and 2. It is possible to ask, from the prediction of TAWSS of Level 3 and knowing the area changes, is it possible to recover the prediction of Levels 1 and 2? The following approach is adopted here. Shear stress is given by
@@
Here, symbols $U$ and $D$ are the characteristic velocity and diameter at any location within the duct. The average velocity at any location can be estimated by knowing the volumetric flow rate and local cross-sectional area. Since we know the total volumetric flow rate and area at any location, $U$ and $D$ can be estimated for local perturbations while $\frac{u*}{z*}$ can be assumed to be unaffected. At a high shear rate, $\mu_{app}$ can be assumed to be a limiting value (Figure 3). Thus, TAWSS can be estimated locally from the characteristic velocity and diameter at every cross-section. These estimates are shown in Figure 14(a) for all viscosity models by black dashed lines.
@@ 3.4 $Q$ criterion
Another critical parameter for studying strain rate and rotational characteristics is the $Q$ criterion. It is based on the velocity gradient tensor, which is given as the summation of a symmetric and anti-symmetric part as follows:
@@
A positive $Q$ value shows the region with a dominance of rotational flow. Similarly, a negative value denotes a zone where rotational flows are not dominant. In Figure 15a, an iso-surface of the $Q$-criterion is shown for $Q$ = 0.1. Similarly, in Figure 15b, it is shown for $Q$ = -0.1. It is clear that the simplification of the artery leads to a uniform flow distribution. Strong vortex structures are created in Level 1 and progressively disintegrate in Level 2 and Level 3.
@@
In the deceleration phase, the $Q$ function swells out, filling the tube volume. During the diastolic flow is weak, and the $Q$ contour created in Level 3 occupies a smaller volume within the tube. Overall, vortex structure weakens considerably as the geometry is simplified. A similar trend is also seen for the contours of $Q$ = -0.1, as shown in Figure 15b. The nature of the distribution of the positive values of the $Q$ function is associated with wall shear stress, whereas negative $Q$ values indicate that strain dominates over rotation, signifying zones of strong deformation rather than vortex formation.
@@ 3.5 Speed Accuracy trade-off
Figure 16: Speed-accuracy analysis is carried out by performing (a) grid independence test for simplified geometries up to 25 slices and comparing their respective friction factors. (b) The maximum number of iterations required to reach the tolerance of 1e-5 in velocity is plotted as a function of the element size. (c) Maximum number of iterations is plotted for the Apostolidis and Beris (A&\&&B) and Newtonian viscosity models for Level 2 (with 25 slices) and Level 3 simplification.
Further, to show the speed increase with geometric simplification, we show the maximum number of iterations required to reach a prescribed convergence criterion as a function of the element size, in Figure 16(b). The data reveal that simplified geometries with 5 to 20 slices require approximately 150 iterations. In contrast, the simplified artery with 25 slices and the geometry with zero slices need around 300 iterations for the smallest element size and 200 iterations for the largest element size. For Level 1 geometry, the number of iterations increases as the element size is decreased, which is consistent with expectations for unstructured meshes over complex three-dimensional geometries. Thus, there is a clear trend of improvement in terms of the number of iterations (and thus, the speed of computations) as we move towards simplified geometric representations. The difference is as large as four times between Level 1 and Level 2 (with 25 slices) at their optimal element sizes.
In Figure 16(c), the Apostolidis and Beris viscosity model is compared with the Newtonian viscosity model with respect to the maximum number of iterations as a function of element size for Level 2 and Level 3 geometries. A significant difference is observed when the viscosity model becomes non-Newtonian. For the converged mesh, nearly 5000 iterations are required for the non-Newtonian viscosity model as compared to the Newtonian. Overall, for meshes with acceptable grid quality, geometric simplification reduces the number of iterations required, nonlinearity of the viscosity model increases the number of iterations, while grid refinement increases the number of iterations.
@@ 5 Conclusions
A detailed study on a patient-specific geometry of the descending aorta with two simplified representations and three viscosity models is conducted. The geometry is extracted from CT scan data shared by a hospital and reconstructed using image processing tools. Realistic time-dependent boundary conditions are applied at the inlet and outlet in terms of pulsatile velocity (mean Re = 313 and peak Re = 1687) and outlet pressure, respectively. For the geometrical approximation, the original artery (termed Level 1) is further simplified into two more levels, denoted as Level 2 and Level 3. Level 2 is based on the cross-sections of Level 1, which is divided into 25 sections, with each part having a representative average diameter. Thus, Level 2 captures some of the undulations and area fluctuations of Level 1 but loses the local asymmetry of the original. Level 3 represents a smooth duct based on the inlet and outlet cross-sectional area of the actual artery. The 3D simulations are performed with three blood rheology models – Newtonian, Carreau-Yasuda, and Apostolidis and Beris. Various aspects of blood flow for all the levels of the descending aorta geometry (with selected blood viscosity models) are studied. The following conclusions are arrived at in the present work.
@@
    The correlation analysis indicates that Level 2 prediction of WSS maintain a strong agreement with Level 1 results, thereby preserving greater accuracy when employing non-Newtonian models. In contrast, this accuracy significantly diminishes in Levels 2 and 3 when using Newtonian viscosity. Pressure remains well-correlated except during the diastolic phases C and D for the Newtonian model.
  4. 4.
@@
    A simple estimate of the actual WSS in the original artery can be obtained by knowing the prediction from Level 3, scaled by the area fluctuation. The estimates thus retrieved are close to the ones obtained with Level 1.
  6. 6.